OpenStudy (anonymous):

the slope-intercept equaion of the line that passes through the point (-2,2) and is parallel to y=4x+7 is? Can someone help me with this PLEASEEE

OpenStudy (anonymous):

Ok, so you need to use Point-Slope form. Basically it will turn out as y-y1=m(x-x1) Find the slope of that line, which is 4. So use the coordinates of that point as your y1 and x1. So y-2=4(x+2) which is y-2=4x+8 which is: y=4x+10

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

No prob haha